# A Critical Look At Tokenwise Reward-Guided Text Generation

### v1

@@ Abstract
Large language models (LLMs) can significantly be improved by aligning to human preferences—the so-called reinforcement learning from human feedback (RLHF). However, the cost of fine-tuning an LLM is prohibitive for many users. Due to their ability to bypass LLM finetuning, tokenwise reward-guided text generation (RGTG) methods have recently been proposed. They use a reward model trained on full sequences to score partial sequences during a tokenwise decoding, in a bid to steer the generation towards sequences with high rewards. However, these methods have so far been only heuristically motivated and poorly analyzed. In this work, we show that reward models trained on full sequences are not compatible with scoring partial sequences. To alleviate this issue, we propose to explicitly train a Bradley-Terry reward model on partial sequences, and autoregressively sample from the implied tokenwise policy during decoding time. We study the property of this reward model and the implied policy. In particular, we show that this policy is proportional to the ratio of two distinct RLHF policies. We show that our simple approach outperforms previous RGTG methods and achieves similar performance as strong offline baselines but without large-scale LLM finetuning.
@@ 1 Introduction
Large language models (LLMs) provide a modern foundation for most if not all text generation tasks [26, 4, 33, 34]. In practice, significant improvements in the quality of text generation are achieved by aligning LLMs to human preferences [32, 24]. This is typically done by reinforcement learning from human feedback (RLHF), which involves two steps: i) learning a reward model from preference data and ii) fine-tuning an LLM to maximize expected rewards by reinforcement learning [44]. Usually, this is done via a reinforcement learning algorithm such as proximal policy optimization [PPO, 29]. Nevertheless, recently, Rafailov et al. [27] showed that the reward modeling step (i) can be bypassed by directly fine-tuning an LLM with preference data, resulting in a method called direct preference optimization (DPO). While this simplifies RLHF, the fine-tuning step (ii) remains prohibitively costly for most users since it requires high-performance computational resources with large GPUs.
@@
In order to alleviate the computational issue above, Khanov et al. [14], Deng and Raffel [9] explored tokenwise reward-guided text generation (RGTG) techniques that avoid any fine-tuning of the LLM. More precisely, the LLM remains frozen (i.e., not finetuned) and the reward model is used at decoding time to directly adjust the softmax scores of the tokens. Unlike DPO, this line of work retains the reward modeling step, but training reward models are typically a much cheaper endeavor than finetuning text-generation LLMs since smaller models can be utilized for reward modeling. Furthermore, reward models are modular: they can easily be composed and reused without any cost to guide text generation in conjunction with any base LLM. In contrast, RLHF via DPO requires fine-tuning of every LLM that we wish to improve based on human preference data.
@@
While RGTG is an interesting alternative to the standard offline RLHF, it is often based on heuristics and still poorly analyzed. For instance, ARGS [14] proposed to simply use a reward model trained on full sequences to score each partial sequence during autoregressive decoding. Meanwhile, Deng and Raffel [9] used a custom tokenwise loss to distill a reward model trained on full sequences. It is thus unclear if these approaches can give rise to a sound tokenwise text generation policy.
@@
In this work, we analyze this common RGTG approach. First, we show that the usage of full-sequence reward models for scoring partial sequences in a tokenwise policy is pathological. To alleviate this, we propose to explicitly train a Bradley-Terry reward model on partial sequences and analyze their property. We prove that this text generation policy is a ratio of two different RLHF policies trained on sequences of different lengths. Ideally, the policy would be derived from a single RLHF policy, but as we shall also show in Section 3, the exact computation of such a policy is intractable. By deriving the sampling policy from a ratio of distinct RLHF policies, we obtain a tractable sampling procedure. This is akin to the argument of Zhao et al. [41] where they use a similar ratio but for deriving a sequential Monte Carlo method.
@@
We empirically validate our analysis on two different text generation datasets on two recent LLMs. Evidence shows that our RGTG approach achieves better performance compared to ARGS, matching the performance of the more expensive, offline PPO and DPO baselines.
In summary, our contributions are:
@@
    We propose to explicitly train a Bradley-Terry reward model on partial sequences and sample from the induced per-token policy induced by it during the decoding time.
@@
    We show that this reward model is a ratio of two distinct RLHF policies over sequences with different lengths. This is a tradeoff that one must make to make tokenwise RGTG free of the pathology and yet still tractable.
@@
    Experiments with recent LLMs such as Llama-2-7b [34] validate our theoretical insights.
@@ 2.1 Reinforcement Learning from Human Feedback
Given a preference dataset $\mathcal{D}=\{({\mathbf{x}}_{k},{\mathbf{y}}_{wk},{\mathbf{y}}_{lk})\}_{k=1}^{K}$ containing $K$ triples of token sequences $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$, Ziegler et al. [44], Ouyang et al. [24] proposed a technique based on reinforcement learning (RL) to align an LLM with the preference dataset. They train a parametric reward model $r_{\phi}({\mathbf{y}}|{\mathbf{x}})$ that assigns a higher score to the “winning” (i.e., preferred) utterance ${\mathbf{y}}_{w}$ and a lower score to the “losing” utterance ${\mathbf{y}}_{l}$. This is done via the Bradley-Terry model [3] which minimizes the loss:
@@
$\max_{\theta}\mathop{\mathbb{E}}_{\begin{subarray}{c}{\mathbf{x}}\sim\mathcal{D},\\ {\mathbf{y}}\sim\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})\end{subarray}}[r_{\phi}({\mathbf{y}}|{\mathbf{x}})]-\beta D_{\mathrm{KL}}[\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})\,\|\,\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})],$ (2)
where the KL term forms a regularizer that ensures that the finetuned model will not differ too much from the reference model. The above optimization problem can be optimized by many RL techniques, including the popular proximal policy optimization (PPO) algorithm [29]. This RL optimization is quite costly in practice due to the size of the LLM.
2.2 Direct Preference Optimization
The above optimization (2) has a closed form solution of the form [25]
@@ 2.2 Direct Preference Optimization
Maximizing the above objective with respect to $\theta$ directly finetunes the LLM without the need to learn a reward model. Furthermore, this maximization is done by supervised learning, which is generally simpler than RL. This approach, known as direct preference optimization [DPO, 27], reduced the cost of RLHF while ensuring that the same finetuned LLM is obtained as RLHF based on PPO. Note, however, that both PPO and DPO-based RLHF are still very costly in practice since they require finetuning (a copy of) the target LLM $\pi_{\text{ref}}$.
@@ 2.3 Reward-Guided Text Generation
In a separate line of work, Khanov et al. [14] proposed reward-guided text generation (RGTG) techniques that do not require any LLM fine-tuning, but can obtain sequences ${\mathbf{y}}$ with high reward. This is done by freezing the reference LLM $\pi_{\text{ref}}$ and at decoding time, the next-token probability $\pi_{\text{ref}}(y^{i}\mid{\mathbf{x}},{\mathbf{y}}^{1:i-1})$ is adjusted by a reward model $r_{\phi}$. More specifically, possible values for $y^{i}$ are scored by a weighted combination of logits of $\pi_{\text{ref}}$ and the rewards :
@@
Note however that it is unclear whether the resulting distribution is equivalent/approximating the RLHF policy in (3). Khanov et al. [14] do train the reward model with the Bradley-Terry loss, but it is trained only with complete sequences, i.e. $r_{\phi}({\mathbf{y}}|{\mathbf{x}})$, while it is used to score partial sequences, i.e. $r_{\phi}({\mathbf{y}}^{1:i}|{\mathbf{x}})$. Hence, it is unclear whether the inferred scores for partial sequences are reasonable. Meanwhile, Deng and Raffel [9] learn the reward model by minimizing a cumulative squared loss to distill a full-sequence reward model instead of using the Bradley-Terry loss (1), making the connection to RLHF policy looser. In Section 3 we show that reward models trained only with complete sequences can assign arbitrary scores to partial sequences and in Section 5 we show empirically that the resulting RGTG policy therefore underperforms that of PPO or DPO.
Along the same lines, Zhao et al. [41] proposed to match each of the marginal distribution of $\pi_{\theta}({\mathbf{y}}^{1:i}|{\mathbf{x}})$ by learning a series of parametric functions $\smash{\{\psi_{\phi_{i}}\}_{i=1}^{|{\mathbf{y}}|}}$. This in turn induces a policy:
@@
The generated sequences ${\mathbf{y}}$ are then approximately equal to the sequences generated by the RLHF policy (3). However, their method is general and does not specifically target RGTG—indeed, Zhao et al. [41] focused on using the implied approximation of the partition function $Z({\mathbf{x}})$.
Finally, Rafailov et al. [28] modifies DPO to obtain a partial-sequence reward model
@@ 3 Pitfalls of RGTG and How to Fix Them
First, we start by analyzing the partial sequence rewards inferred from a reward model trained with full sequences only.
@@
Proof.
Figure 1: A pathology of using a reward model trained on full-sequence to predict partial sequences in decoding-time RGTG. We denote $r^{i}=r(y^{i}|{\mathbf{x}},{\mathbf{y}}^{1:i-1})$. While the total reward over the full sequence ${\mathbf{y}}=(y^{1},\dots,y^{n})$ might be nonzero $c$, it could be in the extreme case that the values over previous partial sequences are all zero—this is a perfectly valid result for a sequence-level reward model (top). This means we can have an unguided decoding, in a reward-guided decoding. By explicitly training $r$ on partial sequences, we could avoid this issue (bottom): While ${\mathbf{y}}$ might achieve the same final reward $c$, nonvanishing reward signals over partial sequences could be avoided.
This leads to an unidentifiability problem—see Fig. 1 for an example. If we learn a reward model based on preferences over full sequences only as proposed by Khanov et al. [14], Deng and Raffel [9], then we may not obtain adequate rewards for partial sequences. As a concrete example, suppose $r$ is a reward model such that (Fig. 1)
@@
This reward model satisfies the identity in (4) and therefore could be the solution when minimizing the Bradley-Terry loss (1). If we use this reward model to sample from the induced RLHF optimal policy in (3), then the token level sampling distribution is the same as for the reference LLM $\pi_{\text{ref}}(y^{i}\mid{\mathbf{x}},{\mathbf{y}}^{1:i-1})$ for all tokens except the last one. This is problematic since RLHF generally changes the token level distribution at each position, not just the last token. Hence the ARGS method [14] may utilize a reward model that does not score partial sequences properly, which in turn may negatively impact token-by-token generation.
Figure 2: The proposed approach to alleviating the problem in Theorem 1. First, we (i) train the reward model $r$ on partial sequences explicitly (when $|{\mathbf{y}}_{w}|\neq|{\mathbf{y}}_{l}|$, we pad to the longest sequence) and (ii) sample from the weighted sum of the logits and the rewards of the next token during decoding. This is in contrast to previous RGTG methods where the reward model is trained on full sequences, but the decoding relies on partial sequence scoring [14, 9].
To alleviate this issue, we propose to explicitly train the reward model with partial sequences—still using the Bradley-Terry model in contrast to Deng and Raffel [9]—as follows (Fig. 2). We create a separate loss function for all prefix lengths $i$:
@@
Then, given that full sequence ${\mathbf{y}}_{w}$ is preferred to full sequence ${\mathbf{y}}_{l}$, we assume that the partial sequence ${\mathbf{y}}_{w}^{1:i}$ is also preferred to the partial sequence ${\mathbf{y}}_{l}^{1:i}$. Strictly speaking, it is hard for human annotators to compare partial sequences due to their incomplete nature and most preference datasets do not include preferences over partial sequences. Nevertheless, we can interpret ${\mathbf{y}}_{w}^{1:i}$ as the prefix of a winning sequence that is preferred over a losing sequence with prefix ${\mathbf{y}}_{l}^{1:i}$. Furthermore, optimizing the partial-sequence objective (5) for all lengths $i$ determines a reward model for all response prefixes that is adequate in the sense that it induces a distribution over partial sequences that approximates the true underlying preference distribution $P^{*}({\mathbf{y}}^{1:i}|{\mathbf{x}})$ instead of assigning arbitrary rewards in the sense of Theorem 1.
This constitutes a tokenwise RGTG method. Algorithm 1 summarizes the decoding procedure. Contrary to the previous approach of Khanov et al. [14], it directly follows the policy induced by the explicitly trained reward model over partial sequences. Meanwhile, compared to Deng and Raffel [9] it leverages the standard Bradley-Terry model instead of a custom squared loss function that distills a full-sequence reward model.
Let us analyze the tokenwise sampling distribution in (6). By the definition of conditional distributions, we can rewrite it as a ratio of two partial sequence distributions:
$\pi(y^{i}|{\mathbf{x}},{\mathbf{y}}^{1:i-1})=\frac{\pi({\mathbf{y}}^{1:i}|{\mathbf{x}})}{\pi({\mathbf{y}}^{1:i-1}|{\mathbf{x}})}$
It is still unclear, however, how does this distribution relates to RLHF policies—the main point of the tokenwise RGTG methods. The following theorem shows how the decoding process by following this distribution relates to RLHF-induced policies.
@@
Proof.
Then, for $i<j$, we have by marginalization:
$=\pi_{\text{ref}}({\mathbf{y}}^{1:i}|{\mathbf{x}})\exp(\beta r({\mathbf{x}},{\mathbf{y}}^{1:i}))\sum_{{\mathbf{y}}^{i+1:j}}\pi_{\text{ref}}({\mathbf{y}}^{i+1:j}|{\mathbf{x}},{\mathbf{y}}^{1:i})\frac{\exp(\beta r({\mathbf{y}}^{1:j}|{\mathbf{x}}))}{\exp(\beta r({\mathbf{y}}^{1:i}|{\mathbf{x}}))}$
∝̸$\not\propto\pi_{\text{RLHF},i}({\mathbf{y}}^{1:i}|{\mathbf{x}}).∝$
$=\frac{\pi_{\text{ref}}({\mathbf{x}},{\mathbf{y}}^{1:i})\exp(\beta r({\mathbf{y}}^{1:i}|{\mathbf{x}}))}{\pi_{\text{ref}}({\mathbf{x}},{\mathbf{y}}^{1:i-1})\exp(\beta r({\mathbf{y}}^{1:i-1}|{\mathbf{x}}))}$ (conditional distribution definition)
This completes the proof of the theorem. ∎
@@
Since all the resulting RLHF decoding policies have merit, one could argue that we can keep things simple by selecting only one policy, perhaps the RLHF policy induced by full sequence preferences (i.e., $\pi_{\text{RLHF}}({\mathbf{y}}|{\mathbf{x}})$). However, as discussed by Rafailov et al. [28], Zhao et al. [41], a conditional distribution over full sequences does not give us an immediate procedure to sample token-by-token. Mathematically, we can derive a token-level policy from a full-sequence policy as follows:
@@
However, the summations in the above equation are exponentially large in the length $|{\mathbf{y}}|$ of the sequences. This exponential complexity was also noted by Zhao et al. [41] who proposed a twisted sequential Monte Carlo technique to approximate computation and mitigate the exponential complexity. In contrast, our approach embraces the multitude of RLHF policies and leverages them in a linear time decoding procedure without any approximation of the partial sequence RLHF policies. The ratio policy (6) described here can thus be seen as a necessary tradeoff if one wants to perform tokenwise RGTG without the pathology in Theorem 1.

### v2

# Acknowledgments and Disclosure of Funding

@@ Abstract
Large language models (LLMs) can be improved by aligning with human preferences through fine-tuning—the so-called reinforcement learning from human feedback (RLHF). However, the cost of fine-tuning an LLM is prohibitive for many users. Due to their ability to bypass LLM fine-tuning, prediction-time tokenwise reward-guided text generation (RGTG) methods have recently been proposed. They use a reward model trained on full sequences to score partial sequences during decoding in a bid to steer the generation towards sequences with high rewards. However, these methods have so far been only heuristically motivated and poorly analyzed. In this work, we show that reward models trained on full sequences are not compatible with scoring partial sequences. To alleviate this issue, we propose to train a Bradley-Terry reward model on partial sequences explicitly, and autoregressively sample from the implied tokenwise policy during decoding time. We study the properties of this reward model and the resulting policy: we show that this policy is proportional to the ratio of two distinct RLHF policies. Our simple approach outperforms previous RGTG methods and performs similarly to strong offline baselines without large-scale LLM finetuning.
@@ 1 Introduction
Large language models (LLMs) provide a modern foundation for most, if not all, text generation tasks (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a; b). In practice, significant improvements in the quality of text generation are achieved by aligning LLMs to human preferences (Stiennon et al., 2020b; Ouyang et al., 2022). This is typically performed via reinforcement learning from human feedback (RLHF), which involves two steps: i) learning a reward model from preference data and ii) fine-tuning an LLM to maximize expected rewards by reinforcement learning (Ziegler et al., 2019b). Usually, this is done via a reinforcement learning algorithm such as proximal policy optimization (PPO, Schulman et al., 2017). Nevertheless, recently, Rafailov et al. (2023) showed that the reward modeling step (i) can be bypassed by directly fine-tuning an LLM with preference data, resulting in a method called direct preference optimization (DPO). While this simplifies RLHF, the fine-tuning step (ii) remains prohibitively costly for most users since it requires high-performance computational resources with large GPUs.
@@
In order to alleviate the computational issue above, Khanov et al. (2024); Deng & Raffel (2023) explored tokenwise reward-guided text generation (RGTG) techniques that avoid any fine-tuning of the LLM. More precisely, the LLM remains frozen (i.e., not finetuned), and the reward model is used at the decoding time to adjust the softmax scores of the tokens directly. Unlike DPO, this line of work retains the reward modeling step, but training reward models are typically a much cheaper endeavor than finetuning text-generation LLMs since smaller models can be utilized for reward modeling. Furthermore, reward models are modular: they can easily be composed and reused without any cost to guide text generation in conjunction with any base LLM. In contrast, RLHF via DPO requires fine-tuning of every LLM that we wish to improve based on human preference data.
@@
While RGTG is an interesting alternative to the standard offline RLHF, it is often based on heuristics and still poorly analyzed. For instance, ARGS (Khanov et al., 2024) proposed to simply use a reward model trained on full sequences to score each partial sequence during autoregressive decoding. Meanwhile, Deng & Raffel (2023) used a custom tokenwise loss to distill a reward model trained on full sequences. Thus, it is unclear if these approaches can give rise to a sound tokenwise text generation policy. Controlled decoding (CD; Mudgal et al., 2024), on the other hand, uses rollouts from the base model along with a reward model trained on full sequences to distill the partial reward.
@@
In this work, we analyze this common RGTG approach. First, we show that the usage of full-sequence reward models for scoring partial sequences in a tokenwise policy is pathological. To alleviate this, we propose to explicitly train a Bradley-Terry reward model on partial sequences. We prove that this text generation policy is a ratio of two different RLHF policies trained on sequences of different lengths. Ideally, the policy would be derived from a single RLHF policy, but as we shall also show in Section 3, the exact computation of such a policy is intractable. By deriving the policy from a ratio of distinct RLHF policies, we obtain a tractable sampling procedure. This is akin to the argument of Zhao et al. (2024) where they use a similar ratio but for deriving a sequential Monte Carlo method.
@@
We empirically validate our analysis on three different text generation datasets on two recent LLMs. Evidence shows that our RGTG approach achieves better performance compared to ARGS and CD, matching the performance of the more expensive, offline PPO and DPO baselines. In summary:
@@
    We thus propose to explicitly train a Bradley-Terry reward model on partial sequences and sample from the induced per-token policy induced by it during the decoding time.
@@
    We show that this reward model induces a ratio of two distinct RLHF policies over sequences with different lengths. This is a tradeoff that one must make to make tokenwise RGTG free of the pathology and yet still tractable.
@@
    Extensive experiments with recent LLMs and various text generation tasks validate our insights.
@@ 2.1 Reinforcement Learning from Human Feedback
Given a preference dataset $\mathcal{D}=\{({\mathbf{x}}_{k},{\mathbf{y}}_{wk},{\mathbf{y}}_{lk})\}_{k=1}^{K}$ containing $K$ triples of token sequences $({\mathbf{x}},{\mathbf{y}}_{w},{\mathbf{y}}_{l})$, Ziegler et al. (2019b) and Ouyang et al. (2022) proposed a technique based on reinforcement learning (RL) to align an LLM with the preference dataset. They train a parametric reward model $r_{\phi}({\mathbf{y}}|{\mathbf{x}})$ that assigns a higher score to the “winning” (i.e., preferred) utterance ${\mathbf{y}}_{w}$ and a lower score to the “losing” utterance ${\mathbf{y}}_{l}$. This is done via the Bradley-Terry model (Bradley & Terry, 1952) which minimizes the loss:
@@
$\max_{\theta}\mathop{\mathbb{E}}_{\begin{subarray}{c}{\mathbf{x}}\sim\mathcal{D},\\ {\mathbf{y}}\sim\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})\end{subarray}}[r_{\phi}({\mathbf{y}}|{\mathbf{x}})]-{\frac{1}{\beta}}D_{\mathrm{KL}}[\pi_{\theta}({\mathbf{y}}|{\mathbf{x}})\,\|\,\pi_{\text{ref}}({\mathbf{y}}|{\mathbf{x}})],$ (2)
where the KL term forms a regularizer that ensures that the finetuned model will not differ too much from the reference model. The above optimization problem can be optimized by many RL techniques, including the popular proximal policy optimization (PPO) algorithm (Schulman et al., 2017). This RL optimization is quite costly in practice due to the size of the LLM.
The above optimization (2) has a closed form solution of the form (Peters & Schaal, 2007)
@@
Maximizing the above objective with respect to $\theta$ directly finetunes the LLM without the need to learn a reward model. Furthermore, this maximization is done by supervised learning, which is generally simpler than RL. This approach, known as direct preference optimization (DPO, Rafailov et al., 2023), reduced the cost of RLHF while ensuring that the same finetuned LLM is obtained as RLHF based on PPO. Note, however, that both PPO and DPO-based RLHF are still very costly in practice since they require finetuning (a copy of) the target LLM $\pi_{\text{ref}}$.
@@ 2.2 Reward-Guided Text Generation
In a separate line of work, Khanov et al. (2024) proposed reward-guided text generation (RGTG) techniques that do not require any LLM fine-tuning, but can obtain sequences ${\mathbf{y}}$ with high reward. This is done by freezing the reference LLM $\pi_{\text{ref}}$ and at decoding time, the next-token probability $\pi_{\text{ref}}(y^{i}\mid{\mathbf{x}},{\mathbf{y}}^{1:i-1})$ is adjusted by a reward model $r_{\phi}$. More specifically, possible values for $y^{i}$ are scored by a weighted combination of logits of $\pi_{\text{ref}}$ and the rewards :
@@
Note however that it is unclear whether the resulting distribution is equivalent/approximating the RLHF policy in (3). Khanov et al. (2024) do train the reward model with the Bradley-Terry loss, but it is trained only with complete sequences, i.e. $r_{\phi}({\mathbf{y}}|{\mathbf{x}})$, while it is used to score partial sequences, i.e. $r_{\phi}({\mathbf{y}}^{1:i}|{\mathbf{x}})$. Hence, it is unclear whether the inferred scores for partial sequences are reasonable. In Section 3 we show that reward models trained only with complete sequences can assign arbitrary scores to partial sequences and in Section 5 we show empirically that the resulting RGTG policy therefore underperforms that of PPO or DPO. Meanwhile, Deng & Raffel (2023) learn the reward model by minimizing a cumulative squared loss to distill a full-sequence reward model instead of using the Bradley-Terry loss (1), making the connection to RLHF policy looser. Mudgal et al. (2024) also distill a partial reward model from a full-sequence reward model, but the tokenwise policy is not the marginal of the full-sequence policy. Nevertheless, tokens are sampled from a different tokenwise RL formulation that follows a similar derivation as RLHF.
Zhao et al. (2024) proposed to match each of the marginal distribution of $\pi_{\theta}({\mathbf{y}}^{1:i}|{\mathbf{x}})$ by learning a series of parametric functions $\smash{\{\psi_{\phi_{i}}\}_{i=1}^{|{\mathbf{y}}|}}$. This in turn induces a policy:
@@
The generated sequences ${\mathbf{y}}$ are then approximately equal to the sequences generated by the RLHF policy (3). However, their method is general and does not specifically target RGTG—indeed, Zhao et al. (2024) focused on using the implied approximation of the partition function $Z({\mathbf{x}})$.
Finally, Rafailov et al. (2024) modifies DPO to obtain a partial-sequence reward model
@@ 3 Pitfalls of RGTG and How to Fix Them
First, we start by analyzing the partial sequence rewards inferred from a reward model trained with full sequences only. Proof in Appendix A.
@@
Figure 1: A pathology of using a reward model trained on full-sequence to predict partial sequences in decoding-time RGTG. We denote $r^{i}=r(y^{i}|{\mathbf{x}},{\mathbf{y}}^{1:i-1})$. While the total reward over the full sequence ${\mathbf{y}}=(y^{1},\dots,y^{n})$ might be nonzero $c$, it could be in the extreme case that the values over previous partial sequences are all zero—this is a perfectly valid result for a sequence-level reward model (top). This means we can have an unguided decoding in a reward-guided decoding. By explicitly training $r$ on partial sequences, we could avoid this issue (bottom): While ${\mathbf{y}}$ might achieve the same final reward $c$, nonvanishing reward signals over partial sequences could be avoided.
This leads to an unidentifiability problem—see Fig. 1 for an example. If we learn a reward model based on preferences over full sequences only as proposed by Khanov et al. (2024) and Deng & Raffel (2023), then we may not obtain adequate rewards for partial sequences. As a concrete example, suppose $r$ is a reward model such that (Fig. 1)
@@
This reward model satisfies the identity in (8) and therefore could be the solution when minimizing the Bradley-Terry loss (1). If we use this reward model to sample from the induced RLHF optimal policy in (3), then the token level sampling distribution is the same as for the reference LLM $\pi_{\text{ref}}(y^{i}\mid{\mathbf{x}},{\mathbf{y}}^{1:i-1})$ for all tokens except the last one. This is problematic since RLHF generally changes the token level distribution at each position, not just the last token. Hence the ARGS method (Khanov et al., 2024) may utilize a reward model that does not score partial sequences properly, which in turn may negatively impact token-by-token generation.
Figure 2: The proposed approach to alleviating the problem in Theorem 1. First, we (i) train the reward model $r$ on partial sequences explicitly (when $|{\mathbf{y}}_{w}|\neq|{\mathbf{y}}_{l}|$, we pad to the longest sequence) and (ii) sample from the weighted sum of the logits and the rewards of the next token during decoding. This is in contrast to previous RGTG methods where the reward model is trained on full sequences, but the decoding relies on partial sequence scoring (Khanov et al., 2024; Deng & Raffel, 2023).
To alleviate this issue, we propose to explicitly train the reward model with partial sequences—still using the Bradley-Terry model in contrast to Deng & Raffel (2023)—as follows (Fig. 2). We create a separate loss function for all prefix lengths $i$:
@@
Then, given that full sequence ${\mathbf{y}}_{w}$ is preferred to full sequence ${\mathbf{y}}_{l}$, we assume that the partial sequence ${\mathbf{y}}_{w}^{1:i}$ is also preferred to the partial sequence ${\mathbf{y}}_{l}^{1:i}$. Strictly speaking, it is hard for human annotators to compare partial sequences due to their incomplete nature and most preference datasets do not include preferences over partial sequences. Nevertheless, we can interpret ${\mathbf{y}}_{w}^{1:i}$ as the prefix of a winning sequence that is preferred over a losing sequence with prefix ${\mathbf{y}}_{l}^{1:i}$. The following lemma shows that the resulting reward model ensures that the probability that a first partial sequence is preferred to a second partial sequence corresponds to the probability that the first sequence is extended to a winning full sequence while the second sequence is extended to a losing full sequence according to the preference data distribution $P_{\text{data}}$. Proof in Appendix A.
Lemma 2.
In the limit of infinite preference data, optimizing a sufficiently expressive reward model according to (4) under the assumption that partial sequences inherit the winning/losing label of full sequences yields a reward model $r_{\phi}$ with the following property:
$\sigma(r_{\phi}({\mathbf{y}}_{1}^{1:i}|{\mathbf{x}})-r_{\phi}({\mathbf{y}}_{2}^{1:j}|{\mathbf{x}}))=P_{\textnormal{data}}({\mathbf{y}}_{1}\succcurlyeq{\mathbf{y}}_{2}|{\mathbf{x}},{\mathbf{y}}_{1}^{1:i},{\mathbf{y}}_{2}^{1:j}),$ (5)
where $P_{\text{data}}$ is the distribution the preference dataset was sampled from and ${\mathbf{y}}_{1}\succcurlyeq{\mathbf{y}}_{2}$ indicates that ${\mathbf{y}}_{1}$ is preferred to ${\mathbf{y}}_{2}$.
Hence, optimizing the partial-sequence objective (4) for all lengths $i$ determines a reward model for all response prefixes that is adequate in the sense that it induces a distribution over partial sequences that approximates the true underlying preference distribution (due to finite data) instead of assigning arbitrary rewards in the sense of Theorem 1.
Algorithm 1 summarizes the decoding procedure. Contrary to the previous approach of Khanov et al. (2024), it directly follows the policy induced by the explicitly trained reward model over partial sequences. Meanwhile, compared to Deng & Raffel (2023) it leverages the standard Bradley-Terry model instead of a custom squared loss function that distills a full-sequence reward model.
Let us now analyze the tokenwise sampling distribution in (6). By the definition of conditional distributions, we can rewrite it as a ratio of two partial sequence distributions: $\pi(y^{i}|{\mathbf{x}},{\mathbf{y}}^{1:i-1})=\nicefrac{{\pi({\mathbf{y}}^{1:i}|{\mathbf{x}})}}{{\pi({\mathbf{y}}^{1:i-1}|{\mathbf{x}})}}$. It is still unclear, however, how does this distribution relates to RLHF policies—the main point of the tokenwise RGTG methods. The following theorem shows how the decoding process by following this distribution relates to RLHF-induced policies. Proof in Appendix A.
@@
Since all the resulting RLHF decoding policies have merit, one could argue that we can keep things simple by selecting only one policy, perhaps the RLHF policy induced by full sequence preferences (i.e., $\pi_{\text{RLHF}}({\mathbf{y}}|{\mathbf{x}})$). However, as discussed by Rafailov et al. (2024) and Zhao et al. (2024), a conditional distribution over full sequences does not give us an immediate procedure for token-wise sampling. Mathematically, we can derive a token-level policy from a full-sequence policy as follows:
@@
However, the summations in the above equation are exponentially large in the length $|{\mathbf{y}}|$ of the sequences. This exponential complexity was also noted by Zhao et al. (2024) who proposed a twisted sequential Monte Carlo technique to approximate computation and mitigate the exponential complexity. In contrast, our approach embraces the multitude of RLHF policies and leverages them in a linear time decoding procedure without any approximation of the partial sequence RLHF policies. The ratio policy (6) described here can thus be seen as a necessary tradeoff if one wants to perform tokenwise RGTG without the pathology in Theorem 1.